OpenStudy (anonymous):

solve the following systems of equations show work please. 2x=5y x+y=1

hero (hero):

y = (2x)/5 y = 1-x Set y = y (2x)/5 = 1-x Solve for x: 5(1-x) = 2x 5 - 5x = 2x 5 = 7x x = 5/7 y = 2/7

hero (hero):

I hope my steps made sense. For these, always solve both equations for y first. Then set y = y, eliminating y, and leaving the equation in terms of x. Then solve for x.

hero (hero):

Thank you flyqueen ;)

OpenStudy (anonymous):

ur welcome show off..lol...just joking i would also showoff if i was as good as u in math..lol ;p

hero (hero):

I don't consider this being good at math.